OpenStudy (anonymous):

Enter in standard form the equation of the line passing through the given point and having the given slope.

OpenStudy (anonymous):

B(2, - 5), m = -2/3

OpenStudy (campbell_st):

the general slope intercept form is y = mx + b m = slope, b = y-intercept so your equation becomes y = -2/3x + b substitute x = 2 and y = -5 to find b does that make sense...

OpenStudy (anonymous):

ya but what would be your final answer?

OpenStudy (campbell_st):

ok so what did you get as the slope intercept answer..?

OpenStudy (campbell_st):

the standard form is Ax + By = C so jut do some manipulation

OpenStudy (anonymous):

sorry still dont get it

OpenStudy (campbell_st):

ok... did you get an equation in slope intercept form...? \[y = -\frac{2}{3} x + b\] so did you substitute x = 2 and y = -5 and find b..?

OpenStudy (anonymous):

okay is b 11/3?

OpenStudy (campbell_st):

close its - 11/3 so the equation is y = -2/3 x - 11/3 mutilpy every term by 3 3y = -2x - 11 just add 2x to both sides of the equation for the answer

OpenStudy (anonymous):

my final answer is y=-2x/3 -11/3

OpenStudy (campbell_st):

thats the slope intercept form of the line your question asked for the standard form Ax + By = C so multiply by 3 3y = -2x - 11 add 2x to both sides of the equation.... to get the standard form.

OpenStudy (anonymous):

what do you meen add 2x on both sides of the equation

OpenStudy (campbell_st):

ok...